OpenStudy (anonymous):

For the expression where x belongs to the set of real numbers find the maximum and minimum value of the following given expression stated as :-

OpenStudy (anonymous):

\[\huge \frac{ 1 }{ \sin^6 x + \cos^6 x }\]

OpenStudy (anonymous):

it simplifies to :- \[\huge \frac{ 1 }{ 1-3\sin^2xcos^2x }\]

OpenStudy (sidsiddhartha):

start with sin^6x+cos^6x

OpenStudy (anonymous):

(sin^2x)^3 + (cos^2x)^3 i did that

OpenStudy (sidsiddhartha):

\[\sin^6x+\cos^6x=(\sin^2x)^3+(\cos^2x)^3\\= (\sin^2 x+ \cos^2x) [ (\sin^2 x)^2 + \sin^2 x \cos^2x +(\cos^2x)^2 ]\]

OpenStudy (sidsiddhartha):

yes?

OpenStudy (anonymous):

(a+b)^3 -3ab(a+b)

OpenStudy (sidsiddhartha):

\[a^3+b^3=(a+b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

(sin^4x + cos^4 x +(sinxcosx)^2)

OpenStudy (sidsiddhartha):

ahh that ab will be -ab

OpenStudy (anonymous):

= -sin^2 x cos ^2x

OpenStudy (sidsiddhartha):

yeash

OpenStudy (sidsiddhartha):

now simplify it more\[\sin^2x=\frac{ 1-\cos2x }{ 2 }\\cos^2x=\frac{ 1+\cos2x }{ 2 }\] put them in the equation

OpenStudy (anonymous):

(cos^2 2x -1)/4

OpenStudy (sidsiddhartha):

so more simplification will give \[\sin^6x+\cos^6x=[(\frac{ 1-\cos2x }{ 2 })^2+(\frac{ 1+\cos2x }{ 2 })^2-\frac{ \sin2x }{ 2 }]\]

OpenStudy (sidsiddhartha):

ok mota bai?

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

ok let me simplify it more, wait

OpenStudy (anonymous):

finee :)

OpenStudy (sidsiddhartha):

yeah now we have \[\sin^6x+\cos^6x=\frac{ 1-(\sin^22x+\sin2x) }{ 2 }\]

OpenStudy (sidsiddhartha):

simplified a little

OpenStudy (sidsiddhartha):

now it will be min when this term \[\sin^2x+\sin2x~will~be~\max\]

OpenStudy (anonymous):

i will show you what i did (a+b)^3 = a^3 +b^3 + 3ab(a+b) a^3 + b^3 = (a+b)^3 -3ab(a+b) so writing it as (sin^2 x ) ^3 + (cos^2 x ) ^3 we get 1- 3(sinxcosx)^2 max of this value how to proceed further

OpenStudy (sidsiddhartha):

\[1-3\sin^2x.\cos^2x\] this will be min when \[\sin^2x.\cos^2x~will~be~\max.\]

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

but what i did is correct right?

OpenStudy (sidsiddhartha):

yes it seems ok

OpenStudy (anonymous):

how to proceed

OpenStudy (sidsiddhartha):

so that term will be max when x=45 deg.

OpenStudy (anonymous):

uhh yes

OpenStudy (anonymous):

got it got it , thank you

OpenStudy (sidsiddhartha):

so 1/4,right?

OpenStudy (anonymous):

yeah so reciprocol of that 4 is answer

OpenStudy (anonymous):

and how to find min

OpenStudy (sidsiddhartha):

same procedure

OpenStudy (anonymous):

min of sin^2x cos^2 x

OpenStudy (sidsiddhartha):

yes

OpenStudy (anonymous):

yeah what would it be

OpenStudy (sidsiddhartha):

aah connection lost again

OpenStudy (sidsiddhartha):

diffrentiate it and see for what value 2nd derivative is positive